# SocialNav: Training Human-Inspired Foundation Model for Socially-Aware Embodied Navigation

### v1

@@ 1 Introduction
As embodied agents become increasingly integrated into everyday social environments, robotic navigation must prioritize not only operational efficiency but also social awareness to ensure safety and compliance with established social norms. However, most existing approaches [singamaneni2024survey, shah2023gnm, shah2023vint, sridhar2024nomad, liu2025citywalker] focus primarily on shortest-path planning and collision avoidance, while overlooking the social compliance essential to real-world deployment (e.g., a robotic guide dog). Consequently, trajectories that appear optimal from a geometric or efficiency perspective may still lead to socially disruptive or inappropriate behaviors, such as jaywalking and traversing restricted zones (e.g., landscaped lawns).
To bridge this gap, we propose SocialNav, a hierarchical foundation model for socially aware navigation that integrates the understanding of social norms with the generation of socially compliant trajectories. SocialNav consists of two core components: (1) the Brain Module, built upon a vision-language model (VLM), which encodes rich social navigation priors and is capable of generating interpretable chain-of-thought (CoT) explanations or explicitly predicting socially traversable regions; and (2) the Action Expert, based on conditional flow matching [lipman2022flow], which translates the semantic priors provided by the Brain Module into robot-executable trajectories that adhere to social norms.
@@
Finally, we introduce the SocNav Benchmark, a high-fidelity evaluation platform blending physics simulation (Isaac Sim [NVIDIA_Isaac_Sim]) and photorealistic rendering (3DGS [kerbl20233d]) in 9 newly captured large-scale social scenes. Our benchmark enables comprehensive comparisons of both fundamental navigation capabilities and social compliance. Our contributions are summarized as follows:
@@ 2.1 Visual Navigation
Visual navigation [bonin2008visual] has evolved from classical SLAM-based methods [cadena2017past, taketomi2017visual, kazerouni2022survey, tsardoulias2016review] to end-to-end learning approaches such as GNM [shah2023gnm], ViNT [shah2023vint], and NoMaD [sridhar2024nomad]. To improve generalization, a line of work has expanded training corpora beyond expert demonstrations. Some methods, such as CityWalker [liu2025citywalker] and MBRA [hirose2025mbra], leverage massive internet-sourced video collections to capture diverse action priors, while others [he2025seeingexperiencingscalingnavigation, cheng2024navila] employ simulation platforms [wu2025urbansim, savva2019habitat] to generate controlled, rule-compliant trajectories at scale. Concurrently, VLMs are used to enhance semantic understanding [rajvanshi2024saynav, cheng2025navila, zhang2024navid, wang2025trackvla, xue2025omninav]. However, these directions still struggle with high-level reasoning and alignment with human values in socially complex scenarios.
@@ 2.2 Social Navigation and Human Preference Alignment
Social navigation requires adherence to social conventions, moving beyond early handcrafted costs [Chen2017, roth2024viplanner] to VLM-based reasoning [lin2025navcot, wang2025think, long2024discuss]. However, VLM reasoning is often disconnected from low-level action generation. Recently, flow matching (FM) [lipman2022flow] has been widely adopted in vision-and-language-action (VLA) models [intelligence2025pi05visionlanguageactionmodelopenworld, yang2025nav] for its ability to model multimodal action distributions, but these models are typically limited to behavior cloning. In the navigation domain, pure imitation learning often lacks the causal understanding needed to robustly adapt to novel or dynamic social situations. Therefore, there is an inherent need for models that can not only generate actions but also align with complex human preferences and social norms. Foundational work, such as GRPO [guo2025deepseek, Shao2024] and Flow-GRPO [liu2505flow], demonstrates combining generative models with online RL for human preference alignment, inspiring our approach.
@@ 3.1 SocNav Dataset
  • Trajectories from internet video ($D_{\text{video}}$). $D_{\text{video}}$, the foundational layer, comprises 2.0 million pseudo-trajectories, meticulously curated from extensive public urban exploration videos to ensure vast visual diversity across global cities, weather conditions, and architectural styles. These raw video streams are transformed into pseudo-trajectories via a scalable pipeline, involving: (1) Dense 3D reconstruction using $\pi^{3}$ [wang2025pi], (2) Metric scale alignment through MoGe [wang2025moge], and (3) Pseudo-trajectory synthesis by sampling diverse point-goal episodes along the reconstructed paths.
  • Trajectories from High-Fidelity Simulated Scenes ($D_{\text{sim}}$). This middle data layer provides 1.7 million diverse trajectories. It comprises over 4,490 high-fidelity 3D scenes and a large-scale dynamic city. Our primary contributions to this layer are: (1) SocialGS, a new dataset of 3,400 real-world scenes (4.5 km²) reconstructed via 3DGS. It diversifies beyond existing residential datasets (e.g., Matterport3D [ramakrishnan2021habitat], InteriorGS [InteriorGS2025]) by covering social environments like shopping malls, streets, and offices. (2) SocCity, a 3.37 km² dynamic urban scene with simulated vehicles and pedestrians in Isaac Sim. Across all scenes, trajectories are generated on manually annotated traversable road networks. These trajectories encompass not only standard on-road navigation but also challenging recovery scenarios, such as near-collisions. This rich dataset enables the model to learn both efficient navigation and robust recovery behaviors.
  • Trajectories from real-world robot data ($D_{\text{real}}$). This layer provides 340K high-quality trajectories collected from autonomous robots deployed in real-world environments (from public datasets including SCAND [karnan2022socially], Huron [hirose2023sacson], Recon [shah2022rapid], and CityWalker teleoperation data [liu2025citywalker]). These trajectories provide ground-truth metric accuracy, physical realism, and sensor consistency, capturing true physical dynamics, sensor noise, and environmental interactions. They are ideal for supervised fine-tuning (SFT) and closing the sim-to-real gap.
@@
  • Navigation Chain-of-Thought (CoT): Leveraging first-person images and curated prompt templates, we prompted Qwen2.5VL-72B [bai2025qwen2] to generate 825K Chain-of-Thought (CoT) samples. These samples, comprising step-by-step textual rationales for navigation decisions, were designed to teach the agent explicit reasoning.
  • General Visual Question Answering (VQA): To ensure the model maintains general-world knowledge, we curated 1 million general VQA samples from [guo2025mammoth, chen2025revisiting, chen2025blip3o, lin2014microsoft, liu2023gres, shao2019objects365]. This task trains the agent to reason about spatial relationships and object properties in the environment.
@@ 4.1 Problem Formulation
We formulate the foundational navigation task as a vision-based, history-conditioned point-goal navigation problem in diverse environments, similar to the setup in CityWalker [liu2025citywalker]. At each time step $t$, the agent receives a sequence of recent monocular visual observations $\bm{O}_{t-n:t}=\{\bm{o}_{t-n},\ldots,\bm{o}_{t}\}$, where $\bm{o}\in\mathbb{R}^{H\times W\times 3}$, along with their associated 2D positional information $\bm{P}_{t-n:t}=\{\bm{p}_{t-n},\ldots,\bm{p}_{t}\}$ with $\bm{p}\in\mathbb{R}^{2}$. Given a specified 2D target location $\bm{g}\in\mathbb{R}^{2}$, the objective is to learn a policy $\pi_{\theta}$ that maps the historical observations and positions to a sequence of future actions:
@@ 4.2 Model Overview
The Action Expert. This action expert specializes in end-to-end trajectory generation. Inspired by recent advancements [black2410pi0, intelligence2025pi05visionlanguageactionmodelopenworld] in action prediction, we leverage conditional flow matching [lipman2022flow] to model a distribution of actions. The module is conditioned on latent semantic features extracted from the last-layer features of the VLM:
@@
Table 1: Open-Loop Evaluation on CityWalker Benchmark [liu2025citywalker]. We evaluate MAOE metric in each critical scenario for all methods. Percentages under scenarios indicate their data proportions. The ”Mean” column shows scenario means averaged over six scenarios; ”All” shows sample means over all data samples.
@@
GNM [shah2023gnm] 16.2 31.1 14.8 12.5 14.7 12.8 11.0 12.1
ViNT [shah2023vint] 16.5 31.1 15.4 12.9 14.8 13.3 11.6 12.6
NoMaD [sridhar2024nomad] 19.1 35.1 18.5 15.6 18.1 14.3 12.8 12.1
CityWalker [liu2025citywalker] 15.2 26.6 14.1 13.9 14.3 12.0 10.4 11.5
@@
GNM* [shah2023gnm] 43.3 62.4 37.0 26.5 28.7
ViNT* [shah2023vint] 45.6 66.2 39.5 31.4 33.8
NoMaD* [sridhar2024nomad] 41.1 60.5 35.4 29.5 31.6
CityWalker [liu2025citywalker] 47.8 64.7 44.7 36.1 36.6
@@ 4.3 Training Methodology
To encourage diverse yet meaningful exploration, we draw inspiration from the flow-based formulation of Flow-GRPO [liu2505flow, wang2025cps], converting the deterministic ordinary differential equation (ODE) of the flow policy into a stochastic differential equation (SDE):
@@ 5 Experiment
To ensure a fair comparison, we evaluate SocialNav in three distinct settings: (1) open-loop benchmark proposed by CityWalker [liu2025citywalker], (2) closed-loop evaluation on our SocNav Benchmark, and (3) real-world robotic deployments.
@@ 5.1 Setup
Metrics. For open-loop evaluation, we use Maximum Average Orientation Error (MAOE) following CityWalker [liu2025citywalker]. Closed-loop performance is measured by success rate (SR; defined as reaching within 3 meters of the goal with fewer than three collisions), route completion (RC), and success weighted by path length (SPL). To assess social compliance, we introduce the Distance Compliance Rate (DCR) and Time Compliance Rate (TCR):
@@
Model Details. We employ Qwen2.5-VL (3B) as the brain module. The action expert is designed as a Diffusion Transformer [peebles2023scalable] with $L=12$ layers, $H=12$ attention heads per layer, and hidden dimension $D=1536$. During inference, the trajectory denoising is performed iteratively for $K=5$ steps.
Training Details. Pre-training Stage: The full model is trained end-to-end using AdamW [loshchilov2017decoupled] (3 epochs, 96 H20 GPUs) with a batch size of 192 and a learning rate of $5\times 10^{-5}$. Finetuning Stage: Only the action expert is fine-tuned on 32 H20 GPUs with a batch size of 256 and a learning rate of $1\times 10^{-5}$. SAFE-GRPO: We further optimize the action expert on 16 H20 GPUs, using a rollout batch size of 128 and a learning rate of $5\times 10^{-7}$.
@@ 5.2 Baselines
We compare SocialNav against state-of-the-art (SOTA) open-source point-based navigation methods. The selected baselines include CityWalker [liu2025citywalker], ViNT [shah2023vint], GNM [shah2023gnm], and NoMaD [sridhar2024nomad]. While ViNT, GNM, and NoMaD were originally designed for image-goal navigation, we re-trained them for point-goal tasks to ensure a fair comparison. In Tab. 2 and Tab. 3, an asterisk (*) denotes the models that were exclusively trained on the $D_{\text{real}}$ dataset.
@@ 5.3 Results on CityWalker Benchmark
We evaluate on the open-loop benchmark introduced by CityWalker [liu2025citywalker]. As shown in Tab. 1, our approach consistently outperforms prior methods across key scenarios, demonstrating improved generalization and robustness. Since ground truth trajectories are human-operated, these results indicate that our method more closely aligns with human social walking norms.
@@ 5.4 Results on our SocNav Benchmark
Navigation Performance.  SocialNav achieves state-of-the-art results across all navigation metrics. Specifically, it significantly outperforms CityWalker [liu2025citywalker], the second-best method, with the concurrent increase of +38.3 in SR, +26.5 in RC and +32.7 in SPL.
@@
GNM* [shah2023gnm] 9/20 10/20 8/20 45.0
ViNT* [shah2023vint] 7/20 12/20 8/20 45.0
NoMaD* [sridhar2024nomad] 9/20 11/20 10/20 50.0
CityWalker [liu2025citywalker] 12/20 13/20 12/20 62.5
@@ 5.5 Results on the Real-World Deployments
We deploy different methods on a cloud server with an NVIDIA A10 GPU to guide a Unitree Go2 robot’s navigation. We evaluate their performance on a set of point-goal navigation tasks across three distinct environments: (1) a street crossing, (2) an office park, and (3) a shopping mall. To ensure a fair comparison, we conduct 20 trials in each environment, using the same starting and ending points for all methods. As shown in Tab. 3, SocialNav achieves a SR of 85, significantly outperforming both NoMaD [sridhar2024nomad] (50) and CityWalker [liu2025citywalker] (62.5). Notably, SocialNav attains 18/20 successful trials in the street crossing environment, reflecting robust behavior in challenging, socially constrained settings. Additionally, SocialNav operates at over 5 Hz during deployment, enabling real-time navigation.
@@ A.4.1 Metric Definitions
  • MAOE. Following CityWalker [liu2025citywalker], we compute the Average Orientation Error (AOE) at each prediction step $k$ as the mean angular difference between the predicted action $\hat{a}_{i_{k}}$ and the ground-truth action $a_{i_{k}}$ over all samples:
@@ B.2 Training Details
Table 9: Open-Loop Evaluation on CityWalker Benchmark [liu2025citywalker]. We evaluate MAOE metric in each critical scenario for all methods. Percentages under scenarios indicate their data proportions. The ”Mean” column shows scenario means averaged over six scenarios; ”All” shows sample means over all data samples.
@@
GNM* [shah2023gnm] 15.2 29.5 13.6 11.9 13.6 12.3 10.4 11.5
ViNT* [shah2023vint] 15.8 30.0 14.6 12.3 14.0 12.7 11.0 12.2
NoMaD* [sridhar2024nomad] 17.8 33.4 16.5 14.8 16.2 13.6 12.2 12.3
CityWalker* [liu2025citywalker] 14.2 25.0 12.8 12.5 13.2 11.5 10.0 11.2

### v2

# arXiv:2511.21135v2[cs.RO] 27 Feb 2026

@@ 1 Introduction
As embodied agents become increasingly integrated into everyday social environments, robotic navigation must prioritize not only operational efficiency but also social awareness to ensure safety and compliance with established social norms. However, most existing approaches [38, 34, 35, 40, 24] focus primarily on shortest-path planning and collision avoidance, while overlooking the social compliance essential to real-world deployment (e.g., a robotic guide dog). Consequently, trajectories that appear optimal from a geometric or efficiency perspective may still lead to socially disruptive or inappropriate behaviors, such as jaywalking and traversing restricted zones (e.g., landscaped lawns).
To bridge this gap, we propose SocialNav, a hierarchical foundation model for socially aware navigation that integrates the understanding of social norms with the generation of socially compliant trajectories. SocialNav consists of two core components: (1) the Brain Module, built upon a vision-language model (VLM), which encodes rich social navigation priors and is capable of generating interpretable chain-of-thought (CoT) explanations or explicitly predicting socially traversable regions; and (2) the Action Expert, based on conditional flow matching [21], which translates the semantic priors provided by the Brain Module into robot-executable trajectories that adhere to social norms.
@@
Finally, we introduce the SocNav Benchmark, a high-fidelity evaluation platform blending physics simulation (Isaac Sim [27]) and photorealistic rendering (3DGS [18]) in 9 newly captured large-scale social scenes. Our benchmark enables comprehensive comparisons of both fundamental navigation capabilities and social compliance. Our contributions are summarized as follows:
@@ 2.1 Visual Navigation
Visual navigation [3] has evolved from classical SLAM-based methods [4, 41, 17, 42] to end-to-end learning approaches such as GNM [34], ViNT [35], and NoMaD [40]. To improve generalization, a line of work has expanded training corpora beyond expert demonstrations. Some methods, such as CityWalker [24] and MBRA [13], leverage massive internet-sourced video collections to capture diverse action priors, while others [12, 8] employ simulation platforms [48, 32] to generate controlled, rule-compliant trajectories at scale. Concurrently, VLMs are used to enhance semantic understanding [29, 9, 51, 45, 49]. However, these directions still struggle with high-level reasoning and alignment with human values in socially complex scenarios.
@@ 2.2 Social Navigation and Human Preference Alignment
Social navigation requires adherence to social conventions, moving beyond early handcrafted costs [7, 31] to VLM-based reasoning [19, 46, 25]. However, VLM reasoning is often disconnected from low-level action generation. Recently, flow matching (FM) [21] has been widely adopted in vision-and-language-action (VLA) models [15, 50] for its ability to model multimodal action distributions, but these models are typically limited to behavior cloning. In the navigation domain, pure imitation learning often lacks the causal understanding needed to robustly adapt to novel or dynamic social situations. Therefore, there is an inherent need for models that can not only generate actions but also align with complex human preferences and social norms. Foundational work, such as GRPO [10, 37] and Flow-GRPO [23], demonstrates combining generative models with online RL for human preference alignment, inspiring our approach.
@@ 3.1 SocNav Dataset
  • Trajectories from internet video ($D_{\text{video}}$). $D_{\text{video}}$, the foundational layer, comprises 2.0 million pseudo-trajectories, meticulously curated from extensive public urban exploration videos to ensure vast visual diversity across global cities, weather conditions, and architectural styles. These raw video streams are transformed into pseudo-trajectories via a scalable pipeline, involving: (1) Dense 3D reconstruction using $\pi^{3}$ [47], (2) Metric scale alignment through MoGe [44], and (3) Pseudo-trajectory synthesis by sampling diverse point-goal episodes along the reconstructed paths.
  • Trajectories from High-Fidelity Simulated Scenes ($D_{\text{sim}}$). This middle data layer provides 1.7 million diverse trajectories. It comprises over 4,490 high-fidelity 3D scenes and a large-scale dynamic city. Our primary contributions to this layer are: (1) SocialGS, a new dataset of 3,400 real-world scenes (4.5 km²) reconstructed via 3DGS. It diversifies beyond existing residential datasets (e.g., Matterport3D [30], InteriorGS [39]) by covering social environments like shopping malls, streets, and offices. (2) SocCity, a 3.37 km² dynamic urban scene with simulated vehicles and pedestrians in Isaac Sim. Across all scenes, trajectories are generated on manually annotated traversable road networks. These trajectories encompass not only standard on-road navigation but also challenging recovery scenarios, such as near-collisions. This rich dataset enables the model to learn both efficient navigation and robust recovery behaviors.
  • Trajectories from real-world robot data ($D_{\text{real}}$). This layer provides 340K high-quality trajectories collected from autonomous robots deployed in real-world environments (from public datasets including SCAND [16], Huron [14], Recon [33], and CityWalker teleoperation data [24]). These trajectories provide ground-truth metric accuracy, physical realism, and sensor consistency, capturing true physical dynamics, sensor noise, and environmental interactions. They are ideal for supervised fine-tuning (SFT) and closing the sim-to-real gap.
@@
  • Navigation Chain-of-Thought (CoT): Leveraging first-person images and curated prompt templates, we prompted Qwen2.5VL-72B [1] to generate 825K Chain-of-Thought (CoT) samples. These samples, comprising step-by-step textual rationales for navigation decisions, were designed to teach the agent explicit reasoning.
  • General Visual Question Answering (VQA): To ensure the model maintains general-world knowledge, we curated 1 million general VQA samples from [11, 5, 6, 20, 22, 36]. This task trains the agent to reason about spatial relationships and object properties in the environment.
@@ 4.1 Problem Formulation
We formulate the foundational navigation task as a vision-based, history-conditioned point-goal navigation problem in diverse environments, similar to the setup in CityWalker [24]. At each time step $t$, the agent receives a sequence of recent monocular visual observations $\bm{O}_{t-n:t}=\{\bm{o}_{t-n},\ldots,\bm{o}_{t}\}$, where $\bm{o}\in\mathbb{R}^{H\times W\times 3}$, along with their associated 2D positional information $\bm{P}_{t-n:t}=\{\bm{p}_{t-n},\ldots,\bm{p}_{t}\}$ with $\bm{p}\in\mathbb{R}^{2}$. Given a specified 2D target location $\bm{g}\in\mathbb{R}^{2}$, the objective is to learn a policy $\pi_{\theta}$ that maps the historical observations and positions to a sequence of future actions:
@@ 4.2 Model Overview
The Action Expert. This action expert specializes in end-to-end trajectory generation. Inspired by recent advancements [2, 15] in action prediction, we leverage conditional flow matching [21] to model a distribution of actions. The module is conditioned on latent semantic features extracted from the last-layer features of the VLM:
@@
Table 1: Open-Loop Evaluation on CityWalker Benchmark [24]. We evaluate MAOE metric in each critical scenario for all methods. Percentages under scenarios indicate their data proportions. The ”Mean” column shows scenario means averaged over six scenarios; ”All” shows sample means over all data samples.
@@
GNM [34] 16.2 31.1 14.8 12.5 14.7 12.8 11.0 12.1
ViNT [35] 16.5 31.1 15.4 12.9 14.8 13.3 11.6 12.6
NoMaD [40] 19.1 35.1 18.5 15.6 18.1 14.3 12.8 12.1
CityWalker [24] 15.2 26.6 14.1 13.9 14.3 12.0 10.4 11.5
@@
GNM* [34] 43.3 62.4 37.0 26.5 28.7
ViNT* [35] 45.6 66.2 39.5 31.4 33.8
NoMaD* [40] 41.1 60.5 35.4 29.5 31.6
CityWalker [24] 47.8 64.7 44.7 36.1 36.6
@@ 4.3 Training Methodology
To encourage diverse yet meaningful exploration, we draw inspiration from the flow-based formulation of Flow-GRPO [23, 43], converting the deterministic ordinary differential equation (ODE) of the flow policy into a stochastic differential equation (SDE):
@@ 5 Experiment
To ensure a fair comparison, we evaluate SocialNav in three distinct settings: (1) open-loop benchmark proposed by CityWalker [24], (2) closed-loop evaluation on our SocNav Benchmark, and (3) real-world robotic deployments.
@@ 5.1 Setup
Metrics. For open-loop evaluation, we use Maximum Average Orientation Error (MAOE) following CityWalker [24]. Closed-loop performance is measured by success rate (SR; defined as reaching within 3 meters of the goal with fewer than three collisions), route completion (RC), and success weighted by path length (SPL). To assess social compliance, we introduce the Distance Compliance Rate (DCR) and Time Compliance Rate (TCR):
@@
Model Details. We employ Qwen2.5-VL (3B) as the brain module. The action expert is designed as a Diffusion Transformer [28] with $L=12$ layers, $H=12$ attention heads per layer, and hidden dimension $D=1536$. During inference, the trajectory denoising is performed iteratively for $K=5$ steps.
Training Details. Pre-training Stage: The full model is trained end-to-end using AdamW [26] (3 epochs, 96 H20 GPUs) with a batch size of 192 and a learning rate of $5\times 10^{-5}$. Finetuning Stage: Only the action expert is fine-tuned on 32 H20 GPUs with a batch size of 256 and a learning rate of $1\times 10^{-5}$. SAFE-GRPO: We further optimize the action expert on 16 H20 GPUs, using a rollout batch size of 128 and a learning rate of $5\times 10^{-7}$.
@@ 5.2 Baselines
We compare SocialNav against state-of-the-art (SOTA) open-source point-based navigation methods. The selected baselines include CityWalker [24], ViNT [35], GNM [34], and NoMaD [40]. While ViNT, GNM, and NoMaD were originally designed for image-goal navigation, we re-trained them for point-goal tasks to ensure a fair comparison. In Tab. 2 and Tab. 3, an asterisk (*) denotes the models that were exclusively trained on the $D_{\text{real}}$ dataset.
@@ 5.3 Results on CityWalker Benchmark
We evaluate on the open-loop benchmark introduced by CityWalker [24]. As shown in Tab. 1, our approach consistently outperforms prior methods across key scenarios, demonstrating improved generalization and robustness. Since ground truth trajectories are human-operated, these results indicate that our method more closely aligns with human social walking norms.
@@ 5.4 Results on our SocNav Benchmark
Navigation Performance.  SocialNav achieves state-of-the-art results across all navigation metrics. Specifically, it significantly outperforms CityWalker [24], the second-best method, with the concurrent increase of +38.3 in SR, +26.5 in RC and +32.7 in SPL.
@@
GNM* [34] 9/20 10/20 8/20 45.0
ViNT* [35] 7/20 12/20 8/20 45.0
NoMaD* [40] 9/20 11/20 10/20 50.0
CityWalker [24] 12/20 13/20 12/20 62.5
@@ 5.5 Results on the Real-World Deployments
We deploy different methods on a cloud server with an NVIDIA A10 GPU to guide a Unitree Go2 robot’s navigation. We evaluate their performance on a set of point-goal navigation tasks across three distinct environments: (1) a street crossing, (2) an office park, and (3) a shopping mall. To ensure a fair comparison, we conduct 20 trials in each environment, using the same starting and ending points for all methods. As shown in Tab. 3, SocialNav achieves a SR of 85, significantly outperforming both NoMaD [40] (50) and CityWalker [24] (62.5). Notably, SocialNav attains 18/20 successful trials in the street crossing environment, reflecting robust behavior in challenging, socially constrained settings. Additionally, SocialNav operates at over 5 Hz during deployment, enabling real-time navigation.
@@ A.4.1 Metric Definitions
  • MAOE. Following CityWalker [24], we compute the Average Orientation Error (AOE) at each prediction step $k$ as the mean angular difference between the predicted action $\hat{a}_{i_{k}}$ and the ground-truth action $a_{i_{k}}$ over all samples:
@@ B.2 Training Details
Table 9: Open-Loop Evaluation on CityWalker Benchmark [24]. We evaluate MAOE metric in each critical scenario for all methods. Percentages under scenarios indicate their data proportions. The ”Mean” column shows scenario means averaged over six scenarios; ”All” shows sample means over all data samples.
@@
GNM* [34] 15.2 29.5 13.6 11.9 13.6 12.3 10.4 11.5
ViNT* [35] 15.8 30.0 14.6 12.3 14.0 12.7 11.0 12.2
NoMaD* [40] 17.8 33.4 16.5 14.8 16.2 13.6 12.2 12.3
CityWalker* [24] 14.2 25.0 12.8 12.5 13.2 11.5 10.0 11.2